Zippy is earning $30,000 per year working for Joes Car Repair. He also has savings of $100,000, on which he is earning 10% annual interest. He decides to leave Joes Car Repair to invest his savings in starting his own car repair business. In the first year, Zippys Speedy Car Repair earns revenues of $200,000 and has explicit costs of $75,000 Zippys economic profit (or loss) in the first year is S. (Round your answer to the nearest dollar.)
Answer #1

Economic profit = Revenue-Total Cost

Total Cost = Explicit Cost+Implicit Cost

Explicit Cost = 75000

Implicit Cost = Salary forgone+ annual interest rate forgone = 30000+10000 = 40000

So, Economic Profit = 200000 - (75000+40000) = 85000

Zippys economic profit in the first year is $85000